What is the equation of the straight line which are equally inclined to the axes on positive side of y-axis?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
slope  = 1, -1  
       Passes through (x1,y1)
(Y-y1)= -(x-x1) or y-y1 = (x-x1)
y-y1= |x-x1| .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)